OpenStudy (anonymous):

A point corresponding to an angle theta in a unit circle what is the csc of theta

OpenStudy (anonymous):

1 / y remember csc = 1 / sin = 1 / y

OpenStudy (anonymous):

For points on the unit circle sin is y/1 cos is x/1 tan is y/x sec is cos flipped upside down csc is sin flipped upside down cot is tan flipped upside down